• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Carlos

    Log entry time 18:40:33 on October 02, 2010

    Entry number 333397

    keyword=Work done during controlled access

    Hisham and I went into the Hall this morning to check BMP signals in the 
    Left arm. Wires 3 and 4 of each BPM seemed to be saturating yesterday 
    with beam.
    
    We checked that attenuators were reading ok (-50mV with a battery of 
    1.5V). Signals were correctly plugged into slot 24 of crate 4. We checked 
    the gate width, it was 60ns wide. We reduced it to 40ns.
    
    In order to troubleshoot with beam, we plugged 4 output channel after the 
    attenuator into the scope (accessible through the web at dvcstlab3):
    
    - Channel 1 of the scope is BPMA, wire 3 (saturating yesterday)
    - Channel 2 : BPMA, wire 4 (saturating yesterday)
    - Channel 3 : BPMB, wire 3 (saturating yesterday)
    - Channel 4 : BPMA, wire 1 (ok yesterday, for reference)